# ICD-LM: Configuring Vision-Language In-Context Demonstrations by Language Modeling

### v2

@@ 4.3.2 Main Result
SITR 66.05 77.69 71.87 83.46 85.05 89.84 93.57 87.98 82.61
SIIR 71.71 84.87 78.29 90.83 93.22 97.80 101.01 95.72 89.91
ICD-LM 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
VQA RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
SITR 40.17 43.58 41.88 46.03 47.5 49.72 50.75 48.50 46.29
SIIR 43.31 47.46 45.39 49.85 50.68 53.23 53.58 51.84 49.69
STTR 44.6 46.75 45.68 47.92 49.05 50.06 49.47 49.13 47.98
ICD-LM 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
4.4 Ablation Studies
We use ablation studies to explore the effects of diverse settings on our approach, including (1) diverse $\mathcal{D}_{\mathcal{M}}$ configurations; (2) diverse scorers $I_{\mathcal{M}}$ in Eq. (3); (3) diverse LM structures; (4) $\mathcal{D}_{\mathcal{M}}$ with longer few-shot ICDs; and (5) randomly ordering the ICD-LM generated ICD sequences.
Diverse $\mathcal{D}_{\mathcal{M}}$ Configurations. We generate 2-shot $\mathcal{D}_{\mathcal{M}}$ by different settings to investigate the corresponding effects. We select three factors for our ablation studies: beam size $b$; the number $n$ of samples in $\mathcal{A}$; and the sampling method of $\mathcal{D}_{S}^{\bm{a}}$, which includes 3 methods: Random: Selecting randomly from $\mathcal{D}_{S}$; Similar Text (Sim-T): Selecting the highest textual similarity sample with anchor sample $\bm{a}$ from $\mathcal{D}_{S}$; Similar Image (Sim-I): Selecting the highest visual similarity sample with anchor sample $\bm{a}$ from $\mathcal{D}_{S}$.
Table 34 show the results for different $\mathcal{D}_{\mathcal{M}}$ configurations on IC and VQA, respectively. We find that ICD-LM can consistently improve the performance compared with the baseline RS The RS means Random Sample ICD retrieval method which is mentioned in Section 4.3.1. in Avg:1$\sim$2. As for length extrapolation capability, only Sim-T gets a lower score than RS. These comparisons confirm the robustness of our method.
Besides, we observe that increasing the beam size has a positive correlation with ICD-LM performance. Specifically, as the beam size increases from $1$ to $5$, the CIDEr score increases by 2.79 for IC and the accuracy increases 1.73 in VQA in Avg:1$\sim$8. This suggests that a diverse $\mathcal{D}_{\mathcal{M}}$ encompasses a broader range of high-quality ICD configurations, which can help train a better ICD-LM. However, an excessively large beam size can negatively impact performance. For instance, in VQA, the accuracy of $b=10$ decays 0.12 than $b=5$ in Avg:1$\sim$8. We hypothesize this drop in performance is due to the introduction of lower-scoring ICD sequences with a large beam size, potentially misleading ICD-LM during training.
Moreover, we observe that using more anchor samples can improve the interpolation performance in both IC and VQA, e.g., when $n$ increases from 1000 to 5000, the Avg:1$\sim$2 CIDEr/accuracy of IC/VQA increases from 83.94/45.39 to 84.32/ 48.75. However, we find that on IC, although the interpolation performance increases when $n$ changes from 3000 to 5000, the extrapolation performance decays, e.g., Avg:3$\sim$8 decrease from 97.60 to 96.52. One possible reason is that using Eq. (3) to build $D_{\mathcal{M}}$ may introduce certain in-domain bias which is beneficial for interpolation while detrimental for extrapolation on IC. Then using more such data may enhance such in-domain bias and thus damage the extrapolation.
For different sample methods of constructing the $\mathcal{D}_{S}^{a}$, we find Random is the best in both IC and VQA. We suppose this is because selecting similar ICDs with the anchor sample from $\mathcal{D}_{S}$ will damage the diversity. Previous study [12] in NLP validates that the diversity of the ICD sequences will also help improve the performance of LLMs.
Table 3: CIDEr of diverse $\mathcal{D}_{\mathcal{M}}$ configurations on IC.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
RS 73.32 82.95 78.1 87.72 93.65 95.81 97.42 93.6 88.5
@@ 4.4 Ablation Studies
Table 4: Accuracy of diverse $\mathcal{D}_{\mathcal{M}}$ configurations on VQA.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
@@
Diverse Scorers $I_{\mathcal{M}}$ for evaluating ICD sequences. To evaluate the quality of ICD configurations, we can also utilize task-specific metrics as $I_{\mathcal{M}}$ to build $D_{\mathcal{M}}$, such as the CIDEr metric in IC. Table 5 compares the results obtained by using prediction confidence and CIDEr as $I_{\mathcal{M}}$.
@@
We can observe that the ICD-LM trained with $\mathcal{D}_{\mathcal{M}}$ using CIDEr scorer can achieve 3.61 higher than Confidence scorer in Avg:1$\sim$2, which means that the CIDEr scorer can assign a more accurate and reasonable score for ICD configurations. However, the length extrapolation capability decreases obviously, which is 3.0 lower than Confidence score in Avg:3$\sim$8. These results evidently show the robustness of Confidence scorer. Although the result shows that the task-specific metric can help the ICD-LM to improve the interpolation ability, it will cost more time and more GPU memory to construct $\mathcal{D}_{\mathcal{M}}$. Specifically, in IC, the time cost of constructing $\mathcal{D}_{\mathcal{M}}$ is approximately 10 times of Confidence.
Diverse LM Structures. In Table  6, we use LSTM [27] as the structure of ICD-LM to generate ICDs for IC and VQA, respectively. We can find that even when using LSTM, our method still achieves excellent performance. In IC, the overall performance improves by 3.64 (92.12 vs. 88.48) compared to the RS baseline, while in VQA, it is improved by 1.38 (49.32 vs. 47.94). However, due to the weak representation learning capability of LSTM, its performance is lower than the Transformer structure,e.g., the scores decrease by 0.33 and 1.99 in IC and VQA, respectively. Overall, these results demonstrate the ability to generate high-quality ICD configurations using various LMs. This validates that employing LM for ICD configuration is a reasonable and effective approach.
@@
ICD-LM(LSTM) 79.73 88.14 83.93 92.36 96.89 96.99 98.63 96.21 92.12
ICD-LM(Transformer) 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
@@
ICD-LM(LSTM) 44.64 48.55 46.60 49.77 50.51 50.71 51.72 50.68 49.32
ICD-LM(Transformer) 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
Table 7: Results of ICD-LM with 4-shot $\mathcal{D}_{\mathcal{M}}$ on IC and VQA.
@@
SITR 66.05 78.66 85.47 86.47 79.16 92.27 95.63 93.95 84.09
SIIR 71.32 84.03 91.38 94.28 85.25 98.63 101.89 100.26 90.25
ICD-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 75.73 85.97 91.28 96.41 87.35 97.56 98.35 97.96 90.88
ICD-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 80.02 88.63 93.41 96.06 89.53 97.26 99.35 98.30 92.45
@@
SITR 39.84 43.60 46.12 47.34 44.23 49.54 50.49 50.02 46.16
SIIR 43.05 47.25 49.68 50.63 47.65 53.16 53.28 53.22 49.51
STTR 43.53 46.89 48.13 48.87 46.86 49.23 49.97 49.60 47.77
ICD-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 44.22 48.11 49.30 51.12 48.19 52.17 52.76 52.47 49.61
ICD-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 46.66 50.83 51.91 52.15 50.39 53.29 53.01 53.15 51.31
Table 8: Results of Random Order of ICD-LM generated ICDs.
2-Shot $\mathcal{D}_{\mathcal{M}}$ 4-Shot $\mathcal{D}_{\mathcal{M}}$
VQA IC VQA IC
Original 50.83 88.63 51.12 85.97
Random Order 50.42 88.56 50.63 85.77
Table 9: Results of Fixed Set ICD-LM on IC and VQA.
@@
Figure 4: Visualizations of using ICDs configured by different methods on VQA. Due to the limited space, we only show the first and the last ICDs. We can find that ICD-LM tends to use more diverse ICDs and thus not lead to short-cut inference.
Longer Few-shot $\mathcal{D}_{\mathcal{M}}$. We further explore the performance of ICD-LM using a 4-shot $\mathcal{D}_{\mathcal{M}}$, as presented in Table 7. It is evident that ICD-LM continues to outperform other retrieval-based methods in Avg:1$\sim$8 metric.
However, we observe a notable performance reduction when the ICD-LM is trained with the 4-shot $\mathcal{D}_{\mathcal{M}}$ compared to the 2-shot $\mathcal{D}_{\mathcal{M}}$. Specifically, for IC, there is a performance decrease of approximately 1.57 in the Avg:1$\sim$8 metric when using the 4-shot $\mathcal{D}_{\mathcal{M}}$ to train ICD-LM. We speculate that this is due to the use of Eq. (2), which may lead to becoming stuck in a local optimum when constructing longer few-shot sequences. Specifically, when selecting the $k$-th ICD, the previous $k-1$ ICDs are held constant. This selection process might not always result in the optimal $k$-shot sequence. When shot number increases, the ICD sequence may further deviate from the global optimum, leading to a decline in ICD-LM performance. Thus leading to poor performance of the ICD-LM generated by this dataset.
Random Order ICD sequence. To validate that our model achieves an effective ICD order, we randomly rearrange the ICD sequences generated by ICD-LM and evaluate the performance. Specifically, we use the ICD-LM trained with 2-shot and 4-shot $\mathcal{D}_{\mathcal{M}}$ and evaluate their performance of the 2-shot and 4-shot ICD configurations, respectively. Table 8 displays our results. It is evident that the original order of ICD configurations generated by the ICD-LM attains the highest score in both VQA and IC, demonstrating that ICD-LM can effectively learn how to order the ICDs.
Fixed Set. During training ICD-LM, we discover that using a high learning rate and not freezing the CLIP model may make ICD-LM converge to a specific solution that for any query input, the ICD configuration is the same and we term this case “Fixed Set”. Table 9 presents the performance of multiple Fixed Set models obtained with different training parameters. The specific training parameter configurations and the generated ICD configurations will be shown in Appendix. We find that the performance of the Fixed Set models fluctuates significantly. For example, in IC, the best version, Fixed Set-2, can outperform the non-Fixed Set by 4.6 points in Avg:1$\sim$8. However, the performance of Fixed Set-2 is poorer than RS in VQA. This phenomenon of Fixed Set demonstrates that for a specific LVLM, there exists a particular ICD configuration that can universally enhance ICL capability. Therefore, we do not need to store a large number of samples for ICD retrieval; instead, we only need to use the Fixed Set samples as the core supporting set. Therefore, it is worthwhile to consider further research in the future to stabilize the performance of the Fixed Set.
5 Conclusion and Limitations
In this paper, motivated from the observation that configuring an ICD sequence is a mirror process of composing a sentence, we propose to build an ICD-LM for configuring ICDs. To achieve this, we construct a dataset containing high-quality ICD sequences to train this ICD-LM. After training, we validate the effectiveness of ICD-LM in accomplishing ICD selection and ordering by comparing it with similarity-based retrieval methods. Also, we deploy extensive ablations to validate the effectiveness of diverse settings of dataset construction and model building strategies.
Supplementary Material
Implementation Details. We provide additional implementation details for each experiment in Table.  10. For all experiments, the batch size is set to 64, the warmup steps to $5\%$ of total training steps, the scheduler as a cosine scheduler, and the optimizer as AdamW [19].
Table 10: Different settings of ICD-LM experiments, where the $n$ is the number of anchor samples in $\mathcal{A}$, $b$ is the beam size, and $l$ is the length of ICD configurations.
Training Parameters $\mathcal{D}_{\mathcal{M}}$ Parameters Avg:1$\sim$8
lr epoch Freeze Adapter $n$ $b$ $l$
[…]
Fixed Set ICD Configurations. We present four ICD configurations of the Fixed Set. Figure 5 displays two ICD configurations for IC, and Figure 6 displays the other two ICD configurations for VQA.

### v4

@@ B.1 Detail Ablation Studies Results On OpenFlamingo
VQA RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
@@
Lever-LM(LSTM) 79.73 88.14 83.93 92.36 96.89 96.99 98.63 96.21 92.12
Lever-LM(Transformer) 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
@@
Lever-LM(LSTM) 44.64 48.55 46.60 49.77 50.51 50.71 51.72 50.68 49.32
Lever-LM(Transformer) 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
Table 10: Results of Lever-LM with 4-shot $\mathcal{D}_{\mathcal{M}}$ on IC and VQA.
@@
SITR 66.05 77.69 83.46 85.05 78.06 89.84 93.57 91.71 82.61
SIIR 71.71 84.87 90.83 93.22 85.16 97.80 101.01 99.40 89.91
Lever-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 75.73 85.97 91.28 96.41 87.35 97.56 98.35 97.96 90.88
Lever-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 80.02 88.63 93.41 96.06 89.53 97.26 99.35 98.30 92.45
@@
SITR 40.17 43.58 46.03 47.5 44.32 49.72 50.75 50.24 46.29
SIIR 43.31 47.46 49.85 50.68 47.83 53.23 53.58 53.41 49.69
STTR 44.6 46.75 47.92 49.05 47.08 50.06 49.47 49.77 47.98
Lever-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 44.55 48.05 50.65 50.98 48.56 52.43 52.92 52.68 49.93
Lever-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 46.66 50.83 51.91 52.15 50.39 53.29 53.01 53.15 51.31
Table 11: Results of Fixed Set Lever-LM on IC and VQA.
@@
B.2 CIDEr Results of Fixed Random ICD Sequence on Image Captioning
To evaluate the effectiveness of Golden-Set, we compare the fixed set learned by Lever-LM with the one constructed by random selection. Specifically, we use OpenFlamingo to randomly select 3 sets of k-shot ICD sequences with different seeds for the image captioning task and then conduct ICL tests. As shown in Table  12, it can be observed that randomly selecting a fixed set of ICD sequences results in relatively poor performance. The Golden-Set outperforms the best Fix Random set (seed 1) in Avg:1$\sim$8 by 16.14. This also demonstrates the importance of high-quality ICD sequences.
Table 12: CIDEr results of a fixed random ICD sequences and Golden-Set on Image Captioning with OpenFlamingo.
Method Avg:1$\sim$2 Avg:3$\sim$8 Avg:1$\sim$8
Random Fix-1 68.95 84.62 79.40
Random Fix-2 64.80 82.15 76.36
Random Fix-3 68.84 86.61 80.68
Golden-Set 91.20 99.63 96.82
B.3 Fine-grained Analysis of Diverse ICL Methods for VQAv2
Table 13 provides a detailed comparison of different ICL methods on the VQAv2 dataset using IDEFICSv1. We compare the accuracy of different question types in VQAv2 and find Lever-LM consistently outperforms other methods, highlighting its effectiveness in VQA tasks.
Table 13: Fine-grained analysis of diverse ICL methods on VQAv2 with IDEFICSv1.
Type Method Avg:1$\sim$2 Avg:4$\sim$8 Avg:1$\sim$8
Yes/No Lever-LM 0.5623 0.5939 0.5833
RS 0.5402 0.5795 0.5664
SIIR 0.5593 0.5853 0.5799
SITR 0.5229 0.5652 0.5511
Other Lever-LM 0.3192 0.3574 0.3446
RS 0.2852 0.3267 0.3129
SIIR 0.2917 0.3349 0.3205
SITR 0.2643 0.3049 0.2914
Counting Lever-LM 0.2782 0.3079 0.2980
RS 0.1323 0.1719 0.1587
SIIR 0.1381 0.1917 0.1738
SITR 0.1051 0.1325 0.1234
B.4 CIDEr Results of Different Lever-LM Sizes in Image Captioning
Table 14 shows the impact of different Lever-LM model sizes on CIDEr scores for image captioning on IDEFICSv1. As shown in Table 14, we evaluate 1-layer/4-layer Transformer decoder layers. We find that the size of Lever-LM has minimal impact on performance. We believe that capturing the ICD sequence distribution is a simple task that can be learned by only a few Transformer Decoder layers. Moreover, our motivation is to use a small model to enhance the ICL performance of a large model, so it is not appropriate to design Lever-LM to be too large.
Table 14: CIDEr results of different Lever-LM sizes in Image Captioning with IDEFICSv1.
Model Size Avg:1$\sim$2 Avg:4$\sim$8 Avg:1$\sim$8
1-layer Transformer (64.2M) 89.48 107.49 101.15
2-layer Transformer (67.4M) 89.58 105.79 100.05
4-layer Transformer (73.7M) 89.12 106.10 100.44
B.5 Lever-LM in NLP domain.
To show Lever-LM is a general method, we train a Lever-LM in a NLP task. Specifically, we use Qwen1.5 [65] 1.8B and generate 2-shot ICD datasets for SST-2 [66], which is a sentiment classification task. The accuracy results are displayed in the Table 15. It can be observed that our Lever-LM outperforms the Random method and STTR in Avg:1$\sim$2 and Avg: 4$\sim$8, demonstrating the potential of Lever-LM in NLP.
Table 15: Accuracy results of diverse ICL methods on SST2 with Qwen1.5-1.8B.
Method Avg:1$\sim$2 Avg:4$\sim$8 Avg:1$\sim$8
RS 0.6227 0.6579 0.6438
STTR 0.6807 0.7312 0.7110
Lever-LM 0.7087 0.7450 0.7305
B.6 Inference Time Comparison between Lever-LM and SIIR.
Table 16 compares the inference time between SIIR and Lever-LM methods on IDEFICSv1. Both methods have similar retrieval times, indicating that Lever-LM’s performance gains do not come at the cost of efficiency.
Table 16: The inference time of different ICL methods with IDEFICSv1.
Method Retrieval Time (s)
SIIR 0.317
Lever-LM 0.328
B.7 Performance on VL-ICL Benchmark Tasks
We test Lever-LM on two tasks of VL-bench [67] due to computation limitation and show the results in Table 17. It can be found Lever-LM achieves higher performance than other retrieval-based methods, validating the generalizability of Lever-LM.
Table 17: Results of diverse methods on two tasks from VL-ICL benchmark with IDEFICSv1.
Task Method Avg:1$\sim$2 Avg:4$\sim$8 Avg:1$\sim$8
VL-ICL CLEVR RS 0.145 0.209 0.188
SIIR 0.170 0.260 0.230
Lever-LM 0.300 0.270 0.280
VL-ICL OCRText RS 0.1923 0.230 0.218
SIIR 0.155 0.164 0.161
Lever-LM 0.262 0.241 0.248
B.8 Performance on IDEFICSv2-8B
We also validate Lever-LM’s generalization ability using IDEFICSv2. IDEFICSv2 is an open-source model specifically designed for ICL. Besides, modern LVLMs with robust ICL ability belong to two mainstream architectures : (1) Flamingo-based (using cross-attention to fuse vision and language, like Open-Flamingo or IDEFICSv1 that are tested in our paper) and (2) LLaVA-based (directly concatenating image and language tokens like IDEFICSv2). Thus, testing with IDEFICSv2 assesses Lever-LM’s generalization across both architectures. Since IDEFICSv2 directly concatenate vision and text tokens, its input sequence will be longer than IDEFICSv1 and 4-shot inference reaches our GPU limitation, thus we report the results of 1-4 shots. The results in Table 18 show that Lever-LM achieves better results than RS and SIIR. Note that RS outperforms SIIR here, one possible reason is that IDEFICSv2 is more likely be damaged by short-cut inference due to similar ICDs are used, while Lever-LM is more robust.
Table 18: CIDEr score of diverse ICL methods on Image Captioning with IDEFICSv2-8B.
Method Avg:1$\sim$2 Avg:3$\sim$4
SIIR 81.93 99.72
RS 87.34 106.07
Lever-LM 100.72 121.68
Appendix C Fixed Set ICD Configurations.
We present four ICD configurations of the Fixed Set. Figure 4 displays two ICD configurations for IC, and Figure 5 displays the other two ICD configurations for VQA.